# Nonlinear anisotropic equilibrium reconstruction in axisymmetric magnetic mirrors

### v3

@@ I Introduction and background
where $\underline{\underline{P}}$ is the pressure tensor and $\vec{B}$ is the magnetic field. We assume an anisotropic pressure tensor with form $\underline{\underline{P}}=p_{\perp}\underline{\underline{I}}-(p_{\perp}-p_{\parallel})\hat{b}\hat{b}$, in which $(\parallel,\perp)$ denote the direction of the pressure relative to the magnetic field direction $\hat{b}=\vec{B}/B$. In the anisotropic case, $p_{\parallel,\perp}$ are functions of $\psi$ and $B$. We exclude flows here as the Alfvénic Mach number $M_{A}^{2}=v^{2}/v_{A}^{2}=\mu_{0}\rho v^{2}/B^{2}\ll 1$, where $v_{A}$ is the Alfvén speed and $\rho$ is the mass density. In WHAM plasmas under consideration here, bias induced shear flows $v\sim 10~\mathrm{km/s}\ll v_{A}\sim 1000~\mathrm{km/s}$. The force balance above may be written in terms of its perpendicular:
@@
In this work, we propose a new method for machine-learning accelerated nonlinear magnetic equilibrium reconstruction in axisymmetric plasmas with anisotropic pressure and high $\beta=2\mu_{0}p/B^{2}$ using a novel kinetic basis function. This yields self-consistent anisotropic pressure and density profiles that allow us to constrain equilibrium reconstructions with a wider array of measurements and provide accurate measures of the plasma $\beta$, total stored energy $W_{\mathrm{tot}}$ and the average ion energy $\langle E_{i}\rangle$, which are key parameters for assessing plasma performance metrics. We describe the Pleiades code used for the reconstructions, the construction of the kinetic basis functions, the machine-learning fitting technique, and then verify our reconstruction technique with synthetic data in Section II. We then apply our equilibrium reconstructions to the Wisconsin High Temperature Superconducting Axisymmetric Mirror (WHAM) Endrizzi et al. (2023) experiments described in Fujii et al.Fujii et al. (2025) in Section III.
@@ II.1 The Pleiades equilibrium solver
$\psi(\vec{x})=\int\mathcal{G}(\vec{x},\vec{x}^{\prime})J_{\phi}(\vec{x}^{\prime})d^{3}x^{\prime},$ (8)
where the Green’s function in cylindrical coordinates is:
@@
where $K$ and $E$ are complete elliptic integrals of the first and second kind respectively and $k^{2}=4RR^{\prime}/[(R+R^{\prime})^{2}+(Z-Z^{\prime})^{2}]$. We discretize equation (8) on an axisymmetric $(R,Z)$ grid. We solve for $\psi$ at each $(R,Z)$ point of interest at $i$ using the $J_{\phi}$ resultant from currents in the plasma at $j$, magnets at $k$, and perfectly conducting boundaries at $l$ at spatial locations $(R^{\prime},Z^{\prime})_{\{j,k,l\}}$. Using this discretization (8) becomes:
@@
To compute an equilibrium, Pleiades evaluates the vacuum values of $\psi$ and $B$, interpolates $p_{\parallel,\perp}(\psi,B)$ onto the the vacuum $\psi$ and $B$ profiles, calculates the plasma contribution to $\psi$ and $B$, re-interpolates $p_{\parallel,\perp}(\psi,B)$ onto the updated $\psi$ and $B$ profiles including the plasma contribution, and then once again recalculates the plasma contributions to $\psi$ and $B$. This process is repeated until the relative change in $\psi$ of the current iteration versus the previous iteration drops below a user input tolerance level. For ease of numerical implementation, we generally represent $p_{\parallel,\perp}(\psi,b)$, where $b=B/B_{0}$ is the magnetic field normalized by the midplane field $B_{0}$.
@@
where $\psi_{n}^{\mathrm{BC}}$ is the flux at $n$ $(R,Z)$ points on the perfectly conducting boundary. Using singular value decomposition (SVD) to obtain a pseudo-inverse of the Green’s function relating currents in the perfect conductor to the surface of the perfect conductor, the above equation can be inverted and solved to obtain a solution for $J_{\phi}^{\mathrm{BC}}$ which approximately enforces the boundary condition. The contribution to the total flux from this boundary condition is then included in the iterative solution to the GS equation and updated on each iteration (as the Green’s function SVD may be precomputed, enforcing this boundary condition is very fast). In WHAM, the vacuum vessel wall is distant from the plasma and the flux loops.Endrizzi et al. (2023) In the limit of a symmetric perfectly conducting wall, Pleiades simulations indicate wall currents have a minimal affect on the diamagnetic flux measurements in present WHAM experiments.
@@ II.2 A semi-analytic kinetic basis for anisotropic mirror pressure profiles
where we have assumed there is some gas dynamically confined cold lumped impurity species with $f_{\mathrm{imp}}=n_{\mathrm{imp}}/n_{\mathrm{GD}}$ and average charge $\langle Z_{\mathrm{imp}}^{2}\rangle$.
@@
To efficiently implement our basis in simulations, a lookup table for $p_{\mathrm{hot,(\parallel,\perp)}}(\psi,b)/\mathcal{A}(\psi)$ and $n_{\mathrm{hot}}(\psi,b)/\mathcal{A}(\psi)$ was constructed for $\theta=45^{\circ}$ and varying values of $Z_{\mathrm{eff}}=1.0$$3.0$ and $T_{e}=0.02$$1.0$ keV (a fixed value of $R_{m}$ was used as there was not significant variation in the solution for $R_{m}=20$100100). A lookup table was preferred to on the fly calculations of $p_{\mathrm{hot,(\parallel,\perp)}}$ and $n_{\mathrm{hot}}$ as such calculations are computationally intensive and would make the equilibrium reconstruction too slow to be used in a routine fashion. In the results presented in this work, 14 terms in the series expansion solution to $f_{\mathrm{hot}}$ in (17) were retained. Additional terms did not improve the solution quality and caused the moment integrals to converge slowly.
@@ II.3 Nonlinear equilibrium reconstruction using Bayesian optimization with Gaussian processes
where $\chi^{2}$ compares $N$ experimental diagnostic measurements $M$ with some Gaussian experimental uncertainty $\sigma_{M}$, to synthetic diagnostic measurements based on the equilibrium reconstruction $\mathcal{M}$. As anisotropic plasmas at high-$\beta$ may have points in the optimization bounds which exceed the $\beta$ limit or are susceptible to instabilities, we utilize a refinement on BOGP, Scalable Constrained Bayesian Optimization (SCBO) Eriksson and Poloczek (2021) based on the Trust Region Bayesian Optimization (TuRBO) technique. Eriksson et al. (2019) The implementation of SCBO used in this work is nearly identical to that described in Eriksson et al.Eriksson and Poloczek (2021) with the exception of the use of Latin-Hypercube sampling McKay et al. (1979) to obtain the initial points rather than Sobol sampling for greater repeatability and somewhat faster convergence.
@@
In addition to using the Gaussian process method as a minimization tool, since we use a $\chi^{2}$ objective function, we can use the Gaussian process surrogate model trained during the minimization to identify the confidence regions in $\chi^{2}$. This allows us to quantify our reconstructions’ uncertainty with very little computational overhead. We identify contours in our surrogate model for $\chi^{2}(\vec{x})$ defining the a confidence interval:
@@ II.4 Verification of Pleiades equilibrium reconstructions using numerical data
Synthetic “measurements” were taken of the excluded magnetic flux at the flux loop locations in the experiment, $(R,Z)=\{(0.25,0.08),\,(0.2,0.35),\,(0.15,0.62)\}~\mathrm{m}$, as well as electron density and temperature at the midplane at six radial points $R\approx\{0.0,\,0.028,\,0.055,\,0.084,\,0.114,\,0.145\}~\textrm{m}$ at the midplane $Z=0~\mathrm{m}$ (corresponding to the radial views of the Thomson system described in Fujii et al.Fujii et al. (2025)). This represents the diagnostic set presently available under nominal WHAM operating conditions. It was assumed in the reconstruction that $T_{e0}(\psi)=T_{e}(\psi,b)$ (no axial variations in $T_{e}$) and that $Z_{\mathrm{eff}}(\psi)=1.0$ (as in the simulation). In these reconstructions, the $T_{e}(\psi)$ and $n_{e}(\psi)$ profiles were set using cubic Hermite interpolation of the synthetic Thomson measurements. The normalized radial shape functions $\mathcal{A}(\psi)/\mathcal{A}_{0}$ and $p_{\mathrm{GD}}(\psi)/p_{\mathrm{GD,0}}$ were assumed to be the same as the normalized Thomson density profile (the normalized electron pressure profile was also considered as a shape function, but the density profile was used as it consistently delivered fits with lower values of $\chi^{2}$). Each measurement from the synthetic diagnostics was assumed to have no uncertainty ($\sigma=0$, analysis of sensitivity to finite diagnostic error on the quality of reconstructions using synthetic data will be the subject of future work), and we used a minimization of modified form of $\chi^{2}=\sum(M_{i}-\mathcal{M}_{i})^{2}/M_{i}^{2}$ to find the normalization of the distribution function $\mathcal{A}_{0}$ on-axis at the midplane and the normalization of the gas dynamic pressure $p_{\mathrm{GD,0}}$ on-axis at the midplane which provided the best fit to the flux loop measurements (as this is not a true $\chi^{2}$ test, we cannot use the uncertainty quantification method currently employed in the code here rigorously, but we plan to develop an enhanced verification suite with realistic noise on measurements in the future).
@@ II.4.1 Gas dynamic Maxwellian
Reconstructions correctly identified the plasma as a gas dynamic Maxwellian, but slightly underestimated the ion energy. This is a result the region around the objective function’s mimimum being very flat leading to some error in the minimum identication at our level of error tolerance in addition to a mismatch between the CQL3D-m ground truth grids and the reconstruction grids (the reconstruction imposes a Gaussian drop near the limiter that slightly reduces average ion energy). Additional diagnostics would improve convergence, and non-uniform profiles would decrease the error resultant from the edge mismatch.
@@ II.4.3 Kinetic sloshing ions
This simulation modeled a WHAM plasma generated by 200 kW of NBI injection. The simulation was initialized with an 100 eV Maxwellian target and evolved to a steady state over 20 ms, sufficient time the majority of the Maxwellian ions to leak out. At the final time-step of the simulation, a reconstruction was performed.
@@ III Reconstructing WHAM experiments
Figure 3: Measurements from the “high-density” 250305121–43 (dark) and “moderate-density” 250306045–64 (light) WHAM experiments. On the left are the Thomson scattering measurements of $n_{e}$ in (a) and $T_{e}$ in (b) as well dashed lines showing their Gaussian fits. On the right are the time traces from high-density shot 250305132 (dark) and moderate-density shot 250306062 (light) showing the interferometer density measurement in (c), the flux loop measurements for flux loops 1 (blue), 2 (orange), and 3 (green) in (d), and forward ECH power (blue) and NBI power (orange) in (e). The black dashed vertical line on the right hand side plots denotes 10.3 ms when the Thomson scattering data was collected.
@@
Figure 4: Plots of the surrogate model for $\chi^{2}$ versus the normalization on the fast ion ion distribution $\mathcal{A_{0}}$ and the Maxwellian ion pressure $p_{\mathrm{GD},0}$ for experiments 250305121–43 (a) and 250306045–64 (b). The joint $\sigma_{1}$ confidence region is shown by the red contour, samples used to construct the model are shown by white dots and the best fit point is shown with a red star. Sloshing ions can be inferred with much greater confidence in the 250306045–64 experiments.
Two sets of experiments utilizing Thomson scattering were analyzed to determine if the presence of sloshing ions could be inferred in WHAM experiments. These were shots 250305121–43 denoted “high density" and shots 250306045–64 denoted “moderate-density" found in Fujii, et al.Fujii et al. (2025) (In these series of shots 250305123, 250305124, 250306046, and 250306050 were excluded as they failed to produce plasmas consistent with the others in their series). Reconstructions were performed at the time of Thomson data collection, $10.3~\mathrm{ms}$, using flux loop measurements that were averaged over the shot series. Measurements from flux loops FL1 and FL2 at $(R,Z)=\{(0.25,0.08),\,(0.2,0.35)\}~\mathrm{m}$ respectively with $\sigma_{M_{i}}=0.1M_{i}$ and the third flux loop FL3 at $(R,Z)=(0.15,0.62)~\mathrm{m}$ with $\sigma_{M_{3}}=0.5M_{3}$ were used in the reconstruction (we used a large $\sigma_{M_{3}}$ as FL3 is prone to significant errors from mechanical noise). The measurements from the diagnostics in these experiments and the Gaussian fits to the Thomson scattering are shown in Figure 3. Following a the same process used in the verification simulations in Section II, the normalized radial shape functions $\mathcal{A}(\psi)/\mathcal{A}_{0}$ and $p_{\mathrm{GD}}(\psi)/p_{\mathrm{GD,0}}$ were assumed to be the same as the normalized Gaussian fits to the Thomson density profiles. These reconstructions used a fixed $Z_{\mathrm{eff}}=2.0$ and $T_{e}(\psi)$ profiles based on Gaussian fits to the Thomson profiles. We then solved for the normalization on the fast-ion profiles $\mathcal{A}_{0}$ and the normalization on the gas dynamic pressure $p_{\mathrm{GD,0}}$ using our equilibrium fitting technique. In the future when more diagnostics are available to constrain the fit and a multiobjective reconstruction algorithm has been implemented, $Z_{\mathrm{eff}}$ and $T_{e}$ will be allowed to be a free parameters. The approach here was chosen as it helps to ensure that the problem is well determined and it performed better than other options in verification tests like those in the previous section. If we do not fix some of the model’s parameters beforehand in this way, we end up in a situation in which we have between 7 and 11 measurements but 14 free parameters (or potentially even more if we chose to use complicated radial shape functions) leaving us with an underdetermined problem.
@@
A clear confounding factor in this analysis is the potential presence of sloshing fast electrons from ECH. In the moderate and high density experiments investigated here, we did not observe strong hard X-ray signals indicative of fast electron generation, but ideally, an additional basis function for the fast electron contribution to the pressure and density profiles would be included in the reconstruction. However, deriving a semi-analytic distribution function basis for the 2D fast electron distribution under quasilinear diffusion is a more complicated problem than the fast ion basis analyzed here. Additionally, at this time WHAM has insufficient diagnostics to meaningfully constrain a three kinetic bases simultaneously without overfitting. A larger number of magnetics measurements, particularly in the regions around the fast electron turning points at $Z\approx\pm(0.7-0.8)~\mathrm{m}$, as well as axially resolved interferometry would improve upon our ability to discern the difference between sloshing ions and electrons in future experiments. We will address both of these issues in future work and WHAM upgrades. However, to make an initial determination as to whether the sloshing ions observed in the moderate-density case were sloshing ions and not sloshing electrons, we devised a test.
@@
here $B_{\mathrm{turn}}=4~\mathrm{T}$ (the fundamental resonance location location for the $110~\mathrm{GHz}$ ECH system in WHAM Endrizzi et al. (2023)), and $A(\psi)$ is a radial shape factor which we constrained based on Thomson scattering using the same method that was used for the fast ions. This basis is ad-hoc, outside of the fact that it satisfies the parallel equilibrium condition (4), but when used to reconstruct equilibria in the SM-1 mirror, it produced excellent agreement the with magnetics signals from ECH experiments.Quon et al. (1985) The test found that the kinetic ion basis function produced better fits to the measured magnetics signals. The fast electron basis was a poor fit in this problem. The $\chi^{2}_{\mathrm{min}}$ solutions when the fast electron basis was used in combination with a Maxwellian basis had $p_{\mathrm{hot}}=0.0$ indicating that the flux loop signals were inconsistent with the presence of sloshing electrons in these plasmas.

### v4

@@ I Introduction and background
where $\underline{\underline{P}}$ is the pressure tensor and $\vec{B}$ is the magnetic field. We assume an anisotropic pressure tensor with form $\underline{\underline{P}}=p_{\perp}\underline{\underline{I}}-(p_{\perp}-p_{\parallel})\hat{b}\hat{b}$, in which $(\parallel,\perp)$ denote the direction of the pressure relative to the magnetic field direction $\hat{b}=\vec{B}/B$. In the anisotropic case, $p_{\parallel,\perp}$ are functions of $\psi$ and $B$. We exclude flows here. The Mach number $M=v^{2}/v_{cs}^{2}\approx v^{2}m_{i}/T_{i}$, where $v$ is the flow velocity, $m_{i}$ is the ion mass, and $T_{i}$ is the ion temperature. The Alfvénic Mach number $M_{A}^{2}=v^{2}/v_{A}^{2}=\mu_{0}\rho v^{2}/B^{2}$, where $v_{A}$ is the Alfvén speed and $\rho$ is the mass density. In the WHAM plasmas under consideration here, perpendicular flows have velocities $v\sim 10~\mathrm{km/s}$, the sound speed $v_{cs}\sim 100~\mathrm{km/s}$, and the Alfvén speed $v_{A}\sim 1000~\mathrm{km/s}$. Therefore, $M_{A}^{2}\ll M^{2}\ll 1$. The value of $M^{2}$ can become appreciable for parallel flows in the expanders and close to the magnetic mirror throat. However, in these regions $M_{A}^{2}$ is often nearly vanishing due to large $B$ and low plasma densities. Thus, there are only very small perturbations to the vacuum fields. The force balance above may be written in terms of its perpendicular:
@@
In this work, we propose a new method for machine-learning accelerated nonlinear magnetic equilibrium reconstruction with automatic uncertainty quantification in axisymmetric plasmas with anisotropic pressure and high $\beta=2\mu_{0}p/B^{2}$ using a novel kinetic basis function. This yields self-consistent anisotropic pressure and density profiles that allow us to constrain equilibrium reconstructions with a wider array of measurements and provide accurate measures of the plasma $\beta$, total stored energy $W_{\mathrm{tot}}$ and the average ion energy $\langle E_{i}\rangle$, which are key parameters for assessing plasma performance metrics. We describe the Pleiades code used for the reconstructions, the construction of the kinetic basis functions, the machine-learning fitting technique, and verify our reconstruction technique with synthetic data in Section II. We then apply our equilibrium reconstructions to the Wisconsin High Temperature Superconducting Axisymmetric Mirror (WHAM) Endrizzi et al. (2023) experiments described in Fujii et al.Fujii et al. (2025) in Section III.
@@ II.1 The Pleiades equilibrium solver
$\psi(\vec{x})=\int_{-\ell}^{\ell}\int_{0}^{a}\mathcal{G}(\vec{x},\vec{x}^{\prime})J_{\phi}(\vec{x}^{\prime})dRdZ,$ (8)
where $\ell$ is the length of the plasma referenced to zero at the midplane, $a$ is the plasma radius, and the Green’s function in cylindrical coordinates is:
@@
where $K$ and $E$ are complete elliptic integrals of the first and second kind respectively and $k^{2}=4RR^{\prime}/[(R+R^{\prime})^{2}+(Z-Z^{\prime})^{2}]$. We discretize (8) on an axisymmetric $(R,Z)$ grid. We solve for $\psi$ at each $(R,Z)$ point of interest at $i$ using the $J_{\phi}$ resultant from currents in the plasma at $j$, magnets at $k$, and perfectly conducting boundaries at $l$ at spatial locations $(R^{\prime},Z^{\prime})_{\{j,k,l\}}$. Using this discretization (8) becomes:
@@
To compute an equilibrium, Pleiades evaluates the vacuum values of $\psi$ and $B$, interpolates $p_{\parallel,\perp}(\psi,B)$ onto the vacuum $\psi$ and $B$ profiles, calculates the plasma contribution to $\psi$ and $B$, re-interpolates $p_{\parallel,\perp}(\psi,B)$ onto the updated $\psi$ and $B$ profiles including the plasma contribution, and then once again recalculates the plasma contributions to $\psi$ and $B$. This process is repeated until the relative change in $\psi$ of the current iteration versus the previous iteration drops below a user input tolerance level. For ease of numerical implementation, we generally represent $p_{\parallel,\perp}(\psi,b)$, where $b=B/B_{0}$ is the magnetic field normalized by the midplane field $B_{0}$.
@@
where $\psi_{n}^{\mathrm{BC}}$ is the flux at $n$ $(R,Z)$ points on the perfectly conducting boundary. Using singular value decomposition (SVD) to obtain a pseudo-inverse of the Green’s function relating currents in the perfect conductor to the surface of the perfect conductor, the above equation can be inverted and solved to obtain a solution for $J_{\phi}^{\mathrm{BC}}$ which approximately enforces the boundary condition. The contribution to the total flux from this boundary condition is then included in the iterative solution to the GS equation and updated on each iteration. As the Green’s function SVD can be precomputed, enforcing this boundary condition is very fast. In WHAM, the vacuum vessel wall is distant from the plasma and the flux loops.Endrizzi et al. (2023) In the limit of a symmetric perfectly conducting wall, Pleiades simulations indicate wall currents have a minimal effect on the diamagnetic flux measurements in present WHAM experiments.
@@ II.2 A semi-analytic kinetic basis for anisotropic mirror pressure profiles
where we have assumed there is some gas dynamically confined cold lumped impurity species with $f_{\mathrm{imp}}=n_{\mathrm{imp}}/n_{\mathrm{GD}}$ and average charge $\langle Z_{\mathrm{imp}}^{2}\rangle$. It is noteworthy that (II.2) implicitly includes the impact of the mirror’s ambipolar potentialPastukhov (1974) despite the potential not appearing explicitly in the MHD equations used to derive the GS equation.
@@
To efficiently implement our basis in simulations, a lookup table for $p_{\mathrm{hot,(\parallel,\perp)}}(\psi,b)/\mathcal{A}(\psi)$ and $n_{\mathrm{hot}}(\psi,b)/\mathcal{A}(\psi)$ was constructed for $\theta=45^{\circ}$ and varying values of $Z_{\mathrm{eff}}=1.0$$3.0$ and $T_{e}=0.02$$1.0$ keV. A fixed value of $R_{m}$ was used as there was not significant variation in the solution for $R_{m}=20$100100. A lookup table was preferred to on the fly calculations of $p_{\mathrm{hot,(\parallel,\perp)}}$ and $n_{\mathrm{hot}}$ as such calculations are computationally intensive and would make the equilibrium reconstruction too slow to be used in a routine fashion. In the results presented in this work, 14 terms in the series expansion solution to $f_{\mathrm{hot}}$ in (17) were retained. Additional terms did not improve the solution quality and caused the moment integrals to converge slowly.
@@ II.3 Nonlinear equilibrium reconstruction using Bayesian optimization with Gaussian processes
where $\chi^{2}$ compares $N$ experimental diagnostic measurements $M$ with some Gaussian experimental uncertainty $\sigma_{M}$, to synthetic diagnostic measurements based on the equilibrium reconstruction $\mathcal{M}$. As anisotropic plasmas at high-$\beta$ may have points in the optimization bounds which exceed the $\beta$ limit or are susceptible to instabilities, we utilize a refinement on BOGP, Scalable Constrained Bayesian Optimization (SCBO) Eriksson and Poloczek (2021) based on the Trust Region Bayesian Optimization (TuRBO) technique. Eriksson et al. (2019) We used BoTorch Balandat et al. (2020) to create an SCBO implementation nearly identical to that described in Eriksson et al.Eriksson and Poloczek (2021) with the exception of the use of Latin-Hypercube sampling McKay et al. (1979) to obtain the initial points rather than Sobol sampling for greater repeatability and somewhat faster convergence.
@@
In addition to using the Gaussian process method as a minimization tool, since we use a $\chi^{2}$ objective function, we can use the Gaussian process surrogate model trained during the minimization to identify the confidence regions in $\chi^{2}$. This allows us to quantify our reconstructions’ uncertainty with very little computational overhead. We identify contours in our surrogate model for $\chi^{2}(\vec{x})$ defining confidence interval:
@@ II.4 Verification of Pleiades equilibrium reconstructions using numerical data
Synthetic “measurements” were taken of the excluded magnetic flux at the flux loop locations in the experiment, $(R,Z)=\{(0.25,0.08),\,(0.2,0.35),\,(0.15,0.62)\}~\mathrm{m}$, as well as electron density and temperature at the midplane at six radial points $R\approx\{0.0,\,0.028,\,0.055,\,0.084,\,0.114,\,0.145\}~\textrm{m}$ at the midplane $Z=0~\mathrm{m}$ (corresponding to the radial views of the Thomson system described in Fujii et al.Fujii et al. (2025)). This represents the diagnostic set presently available under nominal WHAM operating conditions. It was assumed in the reconstruction that $T_{e0}(\psi)=T_{e}(\psi,b)$ (no axial variations in $T_{e}$) and that $Z_{\mathrm{eff}}(\psi)=1.0$ (as in the simulation). In these reconstructions, the $T_{e}(\psi)$ and $n_{e}(\psi)$ profiles were set using cubic Hermite interpolation of the synthetic Thomson measurements. The normalized radial shape functions $\mathcal{A}(\psi)/\mathcal{A}_{0}$ and $p_{\mathrm{GD}}(\psi)/p_{\mathrm{GD,0}}$ were assumed to be the same as the normalized Thomson density profile (the normalized electron pressure profile was also considered as a shape function, but the density profile was used as it consistently delivered fits with lower values of $\chi^{2}$). Each measurement from the synthetic diagnostics was assumed to have no uncertainty ($\sigma=0$). We used a minimization of modified form of $\chi^{2}=\sum(M_{i}-\mathcal{M}_{i})^{2}/M_{i}^{2}$ to find the normalization of the distribution function $\mathcal{A}_{0}$ on-axis at the midplane and the normalization of the gas dynamic pressure $p_{\mathrm{GD,0}}$ on-axis at the midplane which provided the best fit to the flux loop measurements. As this is not a true $\chi^{2}$ test, we could not use the uncertainty quantification method currently employed in the code here, but we plan to develop an enhanced verification suite with realistic noise on measurements in the future.
@@ II.4.1 Gas dynamic Maxwellian
Reconstructions correctly identified the plasma as a gas dynamic Maxwellian, but slightly underestimated the ion energy. This is a result of the region around the objective function’s minimum being very flat leading to some error in the minimum identification at our level of error tolerance in addition to a mismatch between the CQL3D-m ground truth grids and the reconstruction grids (the reconstruction imposes a Gaussian drop near the limiter that slightly reduces average ion energy). Additional diagnostics would improve convergence, and non-uniform profiles would decrease the error resultant from the edge mismatch.
@@ II.4.3 Kinetic sloshing ions
This simulation modeled a WHAM plasma generated by 200 kW of NBI injection. The simulation was initialized with a 100 eV Maxwellian target and evolved to a steady state over 20 ms, sufficient time for the majority of the Maxwellian ions to leak out. At the final time-step of the simulation, a reconstruction was performed.
@@ III Reconstructing WHAM experiments
Figure 3: Measurements from the “high-density” 250305121–43 (dark) and “moderate-density” 250306045–64 (light) WHAM experiments. On the left are the Thomson scattering measurements of $n_{e}$ in (a) and $T_{e}$ in (b) as well as dashed lines showing their Gaussian fits. On the right are the time traces from high-density shot 250305132 (dark) and moderate-density shot 250306062 (light) showing the interferometer density measurement in (c), the flux loop measurements for flux loops 1 (blue), 2 (orange), and 3 (green) in (d), and forward ECH power (blue) and NBI power (orange) in (e). The black dashed vertical line on the right hand side plots denotes 10.3 ms when the Thomson scattering data was collected.
@@
Figure 4: Plots of the surrogate model for $\chi^{2}$ versus the normalization on the fast ion distribution $\mathcal{A_{0}}$ and the Maxwellian ion pressure $p_{\mathrm{GD},0}$ for experiments 250305121–43 (a) and 250306045–64 (b). The joint $\sigma_{1}$ confidence region is shown by the red contour, samples used to construct the model are shown by white dots and the best fit point is shown with a red star. Sloshing ions can be inferred with much greater confidence in the 250306045–64 experiments.
Two sets of experiments utilizing Thomson scattering were analyzed to determine if the presence of sloshing ions could be inferred in WHAM experiments. These were shots 250305121–43 denoted “high-density" and shots 250306045–64 denoted “moderate-density" found in Fujii, et al.Fujii et al. (2025) (In these series of shots 250305123, 250305124, 250306046, and 250306050 were excluded as they failed to produce plasmas consistent with the others in their series). Reconstructions were performed at the time of Thomson data collection, $10.3~\mathrm{ms}$, using flux loop measurements that were averaged over the shot series. Measurements from flux loops FL1 and FL2 at $(R,Z)=\{(0.25,0.08),\,(0.2,0.35)\}~\mathrm{m}$ respectively with $\sigma_{M_{i}}=0.1M_{i}$ and the third flux loop FL3 at $(R,Z)=(0.15,0.62)~\mathrm{m}$ with $\sigma_{M_{3}}=0.5M_{3}$ were used in the reconstruction. We used a large $\sigma_{M_{3}}$ as FL3 is prone to significant errors from mechanical noise. The measurements from the diagnostics in these experiments and the Gaussian fits to the Thomson scattering are shown in Figure 3. Following the same process used in the verification simulations in Section II, the normalized radial shape functions $\mathcal{A}(\psi)/\mathcal{A}_{0}$ and $p_{\mathrm{GD}}(\psi)/p_{\mathrm{GD,0}}$ were assumed to be the same as the normalized Gaussian fits to the Thomson density profiles. These reconstructions used a fixed $Z_{\mathrm{eff}}=2.0$ and $T_{e}(\psi)$ profiles based on Gaussian fits to the Thomson profiles. We then solved for the normalization on the fast-ion profiles $\mathcal{A}_{0}$ and the normalization on the gas dynamic pressure $p_{\mathrm{GD,0}}$ using our equilibrium fitting technique. In the future when more diagnostics are available to constrain the fit and a multiobjective reconstruction algorithm has been implemented, $Z_{\mathrm{eff}}$ and $T_{e}$ will be allowed to be free parameters. The approach here was chosen as it helps to ensure that the problem is well determined and it performed better than other options in verification tests like those in the previous section. If we do not fix some of the model’s parameters beforehand in this way, we end up in a situation in which we have between 7 and 11 measurements but 14 free parameters (or potentially even more if we chose to use complicated radial shape functions) leaving us with an underdetermined problem.
@@
A clear confounding factor in this analysis is the potential presence of sloshing fast electrons from ECH. In the moderate and high density experiments investigated here, we did not observe strong hard X-ray signals indicative of fast electron generation, but ideally, an additional basis function for the fast electron contribution to the pressure and density profiles would be included in the reconstruction. However, deriving a semi-analytic distribution function basis for the 2D fast electron distribution under quasilinear diffusion is a more complicated problem than the fast ion basis analyzed here. Additionally, at this time WHAM has insufficient diagnostics to meaningfully constrain three kinetic bases simultaneously without overfitting. A larger number of magnetics measurements, particularly in the regions around the fast electron turning points at $Z\approx\pm(0.7-0.8)~\mathrm{m}$, as well as axially resolved interferometry would improve upon our ability to discern the difference between sloshing ions and electrons in future experiments. We will address both of these issues in future work and WHAM upgrades. However, to make an initial determination as to whether the sloshing ions observed in the moderate-density case were sloshing ions and not sloshing electrons, we devised a test.
@@
here $B_{\mathrm{turn}}=4~\mathrm{T}$ (the fundamental resonance location for the $110~\mathrm{GHz}$ ECH system in WHAM Endrizzi et al. (2023)), and $A(\psi)$ is a radial shape factor which we constrained based on Thomson scattering using the same method that was used for the fast ions. This basis is ad-hoc, outside of the fact that it satisfies the parallel equilibrium condition (4), but when used to reconstruct equilibria in the SM-1 mirror, it produced excellent agreement with the magnetics signals from ECH experiments.Quon et al. (1985) The test found that the kinetic ion basis function produced better fits to the measured magnetics signals. The fast electron basis was a poor fit in this problem. The $\chi^{2}_{\mathrm{min}}$ solutions when the fast electron basis was used in combination with a Maxwellian basis had $p_{\mathrm{hot}}=0.0$ indicating that the flux loop signals were inconsistent with the presence of sloshing electrons in these plasmas.